You are a racehorse trainer with a horse entered to run in the Japan Cup. Your horse has 60% chance of coming in 1st in the race with a prize of $100,000. You have a 40% chance of losing which means you get nothing in prize money. Your expected value of running your horse in the race is:

 

A. $100,000.

B. $40,000.

C. $60,000.

D. $4,000


1
Expert's answer
2021-11-29T16:26:24-0500

Odds in favour of winning:

"=\frac {Probability Of Winning}{Probability Of Losing}"

"=\frac{\frac{60}{100}}{\frac{40}{100}}"

"=\frac{60}{100}\times\frac{100}{40}=\frac{3}{2}"

"\therefore" the odds in favour of winning are "3:2"

Expected Value:

"=\frac{3}{5}\times100,000=60,000"

=$60,000.